OpenStudy (anonymous):

Amy, Dan, and Mike are competing in a cooking competition. They all used different amounts of milk from a container holding 6 gallons of milk. Amy used 13% of the total milk in the container. Dan used 0.14 of the total milk in the container. Mike used 0.5 gallon of the total milk in the container. Who used the greatest amount of milk from the container? Show your work and explain your answer in words.

OpenStudy (automaticloveletter):

Write 13% in decimal form?

OpenStudy (anonymous):

.13

OpenStudy (danjs):

to change a percent to a decimal, move the decimal left 2 places 13% = .13

OpenStudy (automaticloveletter):

Yeah, now just compare them to find the one that used the most.

OpenStudy (anonymous):

Dan used the most and Mike used the least

OpenStudy (danjs):

i didn't read the gallons word in mike sentence...

OpenStudy (anonymous):

wdym?

OpenStudy (automaticloveletter):

Oh yeah. So do we convert it?

OpenStudy (danjs):

Amy - 13% of 6 gallons = 0.13 * 6 = 0.78 gallons Mike - 0.14 of total = 0.14 * 6 = 0.84 gallons Dan - 0.5 gallons

OpenStudy (danjs):

sorry switch dan and mike name around, mistyped

OpenStudy (danjs):

Amy 0.78 gal Dan 0.84 gal Mike 0.5 gal

OpenStudy (anonymous):

ok ok thank you so much